# Stochastic Approximation with Block Coordinate Optimal Stepsizes

### v1

@@ Acknowledgments
Appendix A PyTorch implementation of BCOS
The following code is also available at https://github.com/facebookresearch/bcos.
1import torch
2from torch.optim import Optimizer
3
4class BCOS(Optimizer):
5 def __init__(self, params, lr, beta=0.9, eps=1e-6,
6 weight_decay=0.1, mode='c', decouple_wd=True):
7
8 defaults = dict(lr=lr, beta=beta, eps=eps, wd=weight_decay)
9 super().__init__(params, defaults)
10
11 if mode not in ['g', 'm', 'c']:
12 raise ValueError(f"BCOS mode {mode} not supported")
13 self.mode = mode
14 self.decouple_wd = decouple_wd # True for BCOSW
15
16 def step(self, closure = None):
17
18 for group in self.param_groups:
19 lr = group["lr"]
20 beta = group["beta"]
21 eps = group["eps"]
22 wd = group["wd"]
23
24 for p in group["params"]:
25 if not p.requires_grad:
26 continue
27
28 state = self.state[p]
29 g = p.grad
30
31 # initialize optimizer states for specific modes
32 if self.mode in ['m', 'c'] and 'm' not in state:
33 state['m'] = g.detach().clone()
34 if self.mode in ['g', 'm'] and 'v' not in state:
35 state['v'] = g.detach().square()
36
37 # decoupled weight decay or absorb in gradient
38 if self.decouple_wd: # p := (1 - lr * wd) * p
39 p.data.mul_(1 - lr * wd)
40 else: # g := g + wd * p
41 g.data.add_(p.data, alpha = wd)
42
43 if self.mode in ['m', 'c']:
44 m = state['m']
45 if self.mode == 'c': # conditional estimator
46 beta_v = 1 - (1 - beta)**2
47 g2 = g.detach().square()
48 v = beta_v * m.square() + (1 - beta_v) * g2
49 # update momentum
50 m.mul_(beta).add_(g.detach(), alpha=1 - beta)
51 d = m
52 else:
53 d = g.detach()
54
55 if self.mode in ['g', 'm']: # EMA estimator
56 v = state['v']
57 v.mul_(beta).add_(d.square(), alpha=1 - beta)
58
59 # BCOS update: p := p - lr * (d / (sqrt(v) + eps))
60 p.data.add_(d.div(v.sqrt() + eps), alpha= - lr)
Appendix B Appendix
B.1 Aiming versus convexity
@@
  • Aiming but not convex: Let $f(x):=\log(x)$ with the optimal solution $x_{\ast}=0$. On the domain of $\mathbf{R}_{+}$, the gradient is $f^{\prime}(x)=\frac{1}{x}$, and thus $\textup{sign}(f^{\prime}(x))=1$ for all $x>0$. Consequently, for any $x\in\mathbf{R}_{+}$, we have
@@
B.2 Proof of Lemma 5.5
Proof.
The proof leverages the Taylor expansion of $\phi(y):=1/\sqrt{y}$. Specifically, we have
@@
where the inequality (69) is a consequence of Assumption B. ∎
B.3 Proof of Lemma 5.6
@@
[…]
@@
Proof.
Let $\delta:=X-\mathbf{E}[X]$ and consider the Taylor expansion of $\phi$ at $\mathbf{E}[X]$:
@@
which, together with the equation $\textup{Cov}(X)=\mathbf{E}[\delta\delta^{T}]$, yields the desired result. ∎
Proof of Lemma 5.6.
We apply Lemma B.1 with $X:=(Y,Z)$ and $\phi(x)=\phi(y,z):=\frac{y}{\sqrt{z}}$. First, the gradient and Hessian of $g$ can be calculated as
@@
$\frac{\partial^{p}\phi}{\partial y^{q}\partial z^{p-q}}=\frac{\partial^{p-q}}{\partial z^{p-q}}\left(\frac{\partial^{q}\phi}{\partial y^{q}}\right)=\begin{cases}\displaystyle 0\quad&\text{if }q\geq 2,\\[8.61108pt] \displaystyle\frac{\partial^{p-1}}{\partial z^{p-1}}\frac{1}{\sqrt{z}}=(-1)^{p-1}\frac{(2p-2)!}{4^{p-1}(p-1)!}z^{-\frac{2p-1}{2}}\quad&\text{if }q=1,\\[8.61108pt] \displaystyle y\cdot\frac{\partial^{p}}{\partial z^{p}}\frac{1}{\sqrt{z}}=(-1)^{p}\frac{(2p)!}{4^{p}p!}yz^{-\frac{2p+1}{2}}\quad&\text{if }q=0.\\ \end{cases}$
@@
$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}\left(\frac{p(2p-2)!}{4^{p-1}(p-1)!}\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}}{\mathbf{E}[Z]^{\frac{2p-1}{2}}}\right]+(-1)^{p}\frac{(2p)!}{4^{p}p!}\mathbf{E}\left[\frac{\mathbf{E}[Y](Z-\mathbf{E}[Z])^{p}}{\mathbf{E}[Z]^{\frac{2p+1}{2}}}\right]\right)$
@@
which is the desired result. ∎
B.4 Bias and variance calculation
@@
  • The conditional estimator of BCOS-c. In this case, we have $d_{t}=m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$ and $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$. Therefore,

### v2

@@ Acknowledgments
Appendix A
A.1 Aiming versus convexity
Our aiming condition in Assumption 1 does not have an upper bound as in the classical condition (32), meaning that it does not require smoothness. Smoothness would give a better bound on the bias and variance of the second-moment estimator of the practical BCOS algorithm, but does not affect stability of the algorithm as long as Assumption 2 holds. Here we focus on the connection of the aiming condition with convexity.
This is a direct consequence of convexity when $x_{\ast}$ is a minimizer of $F$, which is characterized by (e.g., Nesterov04book):
It is easy to come up with examples that satisfy the aiming condition but not convexity. For simplicity, we consider the 1-dimensional case and without loss of generality let $x^{*}=0$. Let $g(x):=\mathbf{E}_{\xi}[g(x,\xi)]$ be any continuous function that satisfies $g(x)>0$ for $x>0$ and $g(x)<0$ for $x<0$, which leads to
$\langle x-x^{*},g(x)\rangle=x\cdot g(x)>0,$
for all $x\neq 0$. For example, consider
$g(x)=x(x-3)^{2}(x+2)^{2}+\lambda x,$
which is a non-monotone polynomial, thus the gradient of a nonconvex polynomial. Yet, this function satisfies the aiming condition (32) for $x^{*}=0$ and for any $\lambda\geq 0$.
However, other than the full-dimensional block case, the aiming condition (67) is not necessarily implied by convexity. As an example, consider the case of single coordinate blocks with uniform SiF, i.e., $\rho_{t,k}=\rho_{t,k^{\prime}}$ for all $k,k^{\prime}\in\{1,\ldots,n\}$. In light of the equivalent formulation using SiF in (34), the corresponding aiming condition becomes
Comparing this condition with convexity, it is easy to construct examples that satisfy one of them while failing the other.
@@ A.1 Aiming versus convexity
  • Aiming but not convex: Let $f(x):=\log(x)$ with the target solution $x_{\ast}=0$. On the domain of $\mathbf{R}_{+}$, the gradient is $f^{\prime}(x)=\frac{1}{x}$, and thus $\textup{sign}(f^{\prime}(x))=1$ for all $x>0$. Consequently, for any $x\in\mathbf{R}_{+}$, we have
@@
A.2 Proof of Lemma 9
Proof The proof leverages the Taylor expansion of $\phi(y):=1/\sqrt{y}$. Specifically, we have
@@ A.2 Proof of Lemma 9
where the inequality (70) is a consequence of Assumption 2.
A.3 Proof of Lemma 10
@@
where $\langle\cdot,\cdot\rangle$ denotes matrix inner product, i.e, $\langle A,B\rangle=\textup{Tr}(A^{T}B)$. For the multi-variable derivatives, we denote $\bm{p}\in\{0,1,2,\ldots\}^{n}$, $|\bm{p}|=p_{1}+\cdots+p_{n}$, $\bm{p}!=p_{1}!\cdots p_{n}!$, and
@@
Proof Let $\delta:=X-\mathbf{E}[X]$ and consider the Taylor expansion of $\phi$ at $\mathbf{E}[X]$:
@@
which, together with the equation $\textup{Cov}(X)=\mathbf{E}[\delta\delta^{T}]$, yields the desired result.
Proof [Proof of Lemma 10] We apply Lemma 15 with $X:=(Y,Z)$ and $\phi(x)=\phi(y,z):=\frac{y}{\sqrt{z}}$. First, the gradient and Hessian of $g$ can be calculated as
$\nabla\phi(x)=\nabla\phi(y,z)=\begin{pmatrix}\displaystyle\frac{1}{z^{1/2}}\\[8.61108pt] \displaystyle-\frac{y}{2z^{3/2}}\end{pmatrix},\qquad\nabla^{2}\phi(x)=\nabla^{2}\phi(y,z)=\begin{bmatrix}\displaystyle 0,&\displaystyle-\frac{1}{2z^{3/2}}\\[8.61108pt] \displaystyle-\frac{1}{2z^{3/2}},&\displaystyle\frac{3y}{4z^{5/2}}\end{bmatrix}$.
@@
$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}(-1)^{p-1}\frac{p(2p-2)!}{4^{p-1}(p-1)!}\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}}{\mathbf{E}[Z]^{\frac{2p-1}{2}}}\right]$
$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}(-1)^{p}\frac{(2p)!}{4^{p}p!}\mathbf{E}\left[\frac{\mathbf{E}[Y](Z-\mathbf{E}[Z])^{p}}{\mathbf{E}[Z]^{\frac{2p+1}{2}}}\right]$
@@
which is the desired result.
A.4 Bias and variance calculation
@@
  • The conditional estimator of BCOS-c. In this case, we have $d_{t}=m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$ and $v_{t}=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}m_{t}+(1-\beta)^{2}g_{t}^{2}$. Therefore, we derive the bias and variance as follows
    $\mathrm{Bias}$ $=|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]|$
    $=\Big|\mathbf{E}_{t}\left[\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)g_{t})+(1-\beta)^{2}g_{t}^{2}\right]$
    $\quad-\mathbf{E}_{t}\left[\left(\beta m_{t-1}+(1-\beta)g_{t}\right)^{2}\right]\Big|$
    $=\Big|\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)\mathbf{E}_{t}[g_{t}])+(1-\beta)^{2}\mathbf{E}_{t}\left[g_{t}^{2}\right]$
    $\quad-\beta^{2}m^{2}_{t-1}-2\beta(1-\beta)m_{t-1}\mathbf{E}_{t}\left[g_{t}\right]-(1-\beta)^{2}\mathbf{E}_{t}\left[g_{t}^{2}\right]\Big|$
    $=\left|2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)\mathbf{E}_{t}[g_{t}]-\mathbf{E}_{t}[g_{t}])\right|$
    $=2\beta^{2}(1-\beta)\left|m_{t-1}\left(m_{t-1}-\mathbf{E}_{t}\left[g_{t}\right]\right)\right|,$
    and
    $\textup{Var}_{t}(v_{t})$ $=\mathbf{E}_{t}\Big[\Big(\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)g_{t})+(1-\beta)^{2}g_{t}^{2}$
    $\quad-\beta^{2}m_{t-1}^{2}-2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)\mathbf{E}_{t}[g_{t}])-(1-\beta)^{2}\mathbf{E}_{t}[g_{t}^{2}]\Big)^{2}\Big]$
    $=\mathbf{E}_{t}\Big[\Big(2\beta(1-\beta)^{2}m_{t-1}(g_{t}-\mathbf{E}_{t}[g_{t}])+(1-\beta)^{2}(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}])\Big)^{2}\Big]$
    $=4\beta^{2}(1-\beta)^{4}m_{t-1}^{2}\mathbf{E}_{t}[(g_{t}-\mathbf{E}_{t}[g_{t}])^{2}]+4\beta(1-\beta)^{4}m_{t-1}\mathbf{E}_{t}[(g_{t}-\mathbf{E}_{t}[g_{t}])(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}])]$
    $\quad+(1-\beta)^{4}\mathbf{E}_{t}\left[\left(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}]\right)^{2}\right]$
    $=4\beta^{2}(1-\beta)^{4}m_{t-1}^{2}\textup{Var}_{t}(g_{t})+4\beta(1-\beta)^{4}m_{t-1}\left(\mathbf{E}_{t}[g_{t}^{3}]-\mathbf{E}_{t}[g_{t}]\mathbf{E}_{t}[g_{t}^{2}]\right)+(1-\beta)^{4}\textup{Var}_{t}\left(g_{t}^{2}\right).$
    The simple alternative $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$ has bias and variance in simpler form: